If the pH of 2.55 L calcium hydroxide solution, Ca(OH)2, is 11.64, calculate the amount (in gram) of calcium hydroxide in the solution.


1
Expert's answer
2022-03-12T01:52:24-0500

pOH=14-pH=-lg[OH^-]


[OH^-]=-10^(-(14-11.64))=4.37*10^(-3) mol/L


Ca(OH)2=Ca^(2+)+2OH^-


[Ca^(2+)]=4.37*10^(-3)/2=2.18*10^(-3) mol/L


mol Ca^(2+)=2.55*2.18*10^(-3)=5.57*10^(-3)


m=5.57*10^(-3)mol*40.078g/mol=0.223g